Calculate the density of oxygen, O2, under each of the following conditions:
STP
1.00 bar and 15.0 ∘C
1
Expert's answer
2020-10-14T05:44:35-0400

The best way to figure this out is to work with PV=nRT first, then put in the numbers.


PV=nRT and n=g/MW (g=# of grams and MW =molecular weight of gas)

now

PV=(g/MW)*RT and V is in liters so solve for g/l

[P*(MW)]/RT = g/V (or g/l)


now plug in the numbers that you have:

at STP: T = 273 K anf P=1 atm

[1*32]/(0.08206*273) = 1.43 g/l


at 1bar=1 atm and 15 deg C (288 K):

[1*32]/(0.08206*288) = 1.35 g/l


This makes sense since as temperature decreases, gases become more dense.